OpenStudy (unklerhaukus):

Evaluate the sum at \(x=\tfrac\pi2\) \[S(x) =\frac{1}2+\frac2\pi\sum\limits_{n=1,3,5...}^\infty\frac{\sin({n x})}n\\=\frac{1}2+\frac2\pi\sum\limits_{r=0}^\infty\frac{\sin({(2r-1) x})}{2r-1}\\ \]

OpenStudy (unklerhaukus):

\[S\left(\tfrac\pi2\right) =\frac{1}2+\frac2\pi\sum\limits_{r=0}^\infty\frac{\sin(\pi r-\tfrac\pi2)}{2r-1}\\ =\frac{1}2+\frac2\pi\sum\limits_{r=0}^\infty\frac{(-1)^{r-1}}{2r-1}\]

OpenStudy (anonymous):

looks to me a series of log(1+x) and log(1-x)

OpenStudy (unklerhaukus):

how

OpenStudy (anonymous):

also r should start in 1 and not 0

OpenStudy (unklerhaukus):

hmm

OpenStudy (anonymous):

see this http://www.wolframalpha.com/input/?i=series+log%281%2Bx%29

OpenStudy (unklerhaukus):

but that has even terms in it

OpenStudy (anonymous):

eliminate using series for log (1-x) try it if it not works use ix instead of x

OpenStudy (unklerhaukus):

im not sure how to do that

OpenStudy (anonymous):

see that \[\frac{1}{2}\frac{\log(1+x)}{\log(1-x)}=x+\frac{x^3}{3}+\frac{x^5}{5}...\] Replace x by ix

OpenStudy (anonymous):

now do you see it

OpenStudy (anonymous):

Signs will now alternate

OpenStudy (unklerhaukus):

is there another way to do this?

OpenStudy (unklerhaukus):

@mukushla

OpenStudy (experimentx):